OpenStudy (anonymous):

the Martians ask you to explain one last thing, Ultimate Math Ambassador. Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will always result in the same number. You will use the inverse function that you created in problem number 5 for g(x

OpenStudy (anonymous):

i need help how to solve

OpenStudy (anonymous):

Function: f(x) = 4x+8 Inverse: g(x) = x+8 / 4

OpenStudy (anonymous):

@satellite73 im not trying to cheat or get answers i just need someone to show me and explain if you could please

OpenStudy (anonymous):

if you really want english, you could say: the first function \(f\) says take a number, a) multiply it by 4 then b) add 8 if you do the reverse operations in the reverse order, you get back to what you started with the reverse operations in the reverse order would be a) subtract 8 b) divide by 4, i.e. \(f^{-1}(x)=\frac{x-8}{4}\)

OpenStudy (anonymous):

which, by the way, means your inverse \(g(x)\) is wrong

OpenStudy (anonymous):

i am assuming in problem 5, which i do not see, that you were supposed to find the inverse function of \(f(x)=4x+8\) which is \(f^{-1}(x)=\frac{x-8}{4}\)

OpenStudy (anonymous):

or as you called it "\(g(x)=\frac{x-8}{4}\)"

OpenStudy (anonymous):

? im confused how would it be wrong i did it the way you said first and my instructor had me resub mit it as the way im showing you

OpenStudy (anonymous):

here is number 5. f(x) = 4x + 8 y = 4x + 8 x = 4y + 8 +8 +8 x + 8 4y ---- = ---- = y 4 4y x+8 ----- = (f-1(x) 4

OpenStudy (anonymous):

i don't see problem 5 does it ask you to find the inverse of \(f(x)=4x+8\) ?

OpenStudy (anonymous):

you have a mistake in your calculation you need so SUBTRACT 8 from both is sides, not add

OpenStudy (anonymous):

\[y=4x+8\\ x=4y+8\\ x-8=4y\\ \frac{x-8}{4}=y\]

OpenStudy (anonymous):

the lines you wrote here x = 4y + 8 +8 +8 are not right you want to get \(y\) by itself, you have to subtract not add i.e. x = 4y + 8 -8 -8 x-8=4y

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

my teacher said it that was the correct way

OpenStudy (anonymous):

one of two things is possible 1) you misunderstood your teacher 2) your teacher was mistaken i won't try to guess which

OpenStudy (anonymous):

ok so how do i do this the Martians ask you to explain one last thing, Ultimate Math Ambassador. Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will always result in the same number. You will use the inverse function that you created in problem number 5 for g(x

OpenStudy (anonymous):

that is the question i answered above

OpenStudy (anonymous):

ik but i mean liike can you do it again but explain so i can do it on my own

OpenStudy (anonymous):

i cannot say it in other words read my first post, see if it makes sense to you if i write it again it will be the same

OpenStudy (anonymous):

multiply by for in g(x) = x+8 / 4 or f(x) = 4x+8

OpenStudy (anonymous):

by 4 *